OpenStudy (anonymous):

Find a^3 + b^3 if a+b=5 & a * b=2

OpenStudy (freckles):

cube both sides of a+b=5 then expand and factor a bit

Directrix (directrix):

Hint from @anwar a+b =5 Cube both sides (a+b)^3 =5^3 which gives a^3 + 3a^2b + 3ab^2 + b^3 = a^3 +b^3 +3ab*(a+b) = 125 put the values of ab and a+b and solve for the value of a^3 + b^3.

Directrix (directrix):

Post what you get, okay? @Cherryboa